# Transferable Post-training via Inverse Value Learning

### v1

@@ Abstract
As post-training processes utilize increasingly large datasets and base models continue to grow in size, the computational demands and implementation challenges of existing algorithms are escalating significantly. In this paper, we propose modeling the changes at the logits level during post-training using a separate neural network (i.e., the value network). After training this network on a small base model using demonstrations, this network can be seamlessly integrated with other pre-trained models during inference, enables them to achieve similar capability enhancements. We systematically investigate the best practices for this paradigm in terms of pre-training weights and connection schemes. We demonstrate that the resulting value network has broad transferability across pre-trained models of different parameter sizes within the same family, models undergoing continuous pre-training within the same family, and models with different vocabularies across families. In certain cases, it can achieve performance comparable to full-parameter fine-tuning. Furthermore, we explore methods to enhance the transferability of the value model and prevent overfitting to the base model used during training. [^1] Our code will be open-source at https://github.com/luxinyu1/inverse-value-learning
@@
Xinyu Lu1,2, Xueru Wen1,2 Yaojie Lu1, Bowen Yu3, Hongyu Lin1, Haiyang Yu3, Le Sun1, Xianpei Han1, Yongbin Li3 1Chinese Information Processing Laboratory Institute of Software, Chinese Academy of Sciences 2University of Chinese Academy of Sciences 3Alibaba Group {luxinyu2021,wenxueru2022,luyaojie,hongyu,sunle,xianpei}@iscas.ac.cn {yubowen.ybw,yifei.yhy,shuide.lyb}@alibaba-inc.com
@@ 1 Introduction
We systematically investigate practical implementations for this framework. Specifically, we first study the importance of pre-trained weights in modeling the residual logits. Then, we examine two architectures for integrating the value network with pre-trained models: Cascade and Residual connections. Our experiments reveal that the residual connection scheme, where the value network predicts residuals based solely on previous text inputs, demonstrates superior transferability and efficiency. To improve the generalization ability of the value network, we incorporate regularization techniques, such as norm constraints, to mitigate overfitting to the base model during training. Additionally, we introduce a vocabulary mapping algorithm to facilitate effective cross-vocabulary transfer.
We assess our approach on multiple datasets and tasks, including general instruction following, zero-shot capabilities, and few-shot learning scenarios. Our results show that the value network exhibits broad transferability across pre-trained models of varying parameter sizes within the same model family, models undergoing continual pre-training, and even across different model families. In certain cases, our framework reaches close performance with heavy full-parameter finetuning. This highlights the potential of our method for efficient and practical application.
@@ 2 Problem Statement and Background
The post-training of LLMs has become increasingly important for enhancing model performance and adapting to specific tasks (Dubey et al., 2024). However, as model sizes continue to grow, the computational and time costs of training the entire model have also increased significantly. In some cases, updating all the weights of the model is impractical, prompting researchers to explore alternative methods, such as updating additional low-rank parameters (Hu et al., 2022) or interfering the representations (Wu et al., 2024) while freezing all the original pre-trained weights.
The inconsistency of the parameter space and representation space between different models stands as a key challenge for transferable post-training. This is due to the differences in model structure, hidden state dimensions, etc. In contrast, the logits space has better sharing properties and can serve as a communication channel for model interaction, enabling the transfer of capabilities between models (Hinton, 1999). Specifically, the impact of post-training on the logits of a pretrained model can be formally expressed as:
@@
where $\log p_{\Delta}$ represents the change in the logits space during post-training. A series of studies have focused on exploring $\log p_{\Delta}$. Mitchell et al. (2023) and Liu et al. (2024) obtained $\log p_{\Delta}$ by comparing an existing instruction-following model with its corresponding base pre-trained model, i.e., $\log p_{\Delta}(y_{t}|{\bm{x}},{\bm{y}}_{<t})=\log\dfrac{p_{\pi^{*}}(y_{t}|{\bm{x}},{\bm{y}}_{<t})}{p_{\text{base}}(y_{t}|{\bm{x}},{\bm{y}}_{<t})}$. This difference can be regarded as an advantage function (Mitchell et al., 2023), or an implicit reward signal (Rafailov et al., 2024). Furthermore,  Khanov et al. (2024) directly uses predefined reward functions as $\log p_{\Delta}$ to guide the Top-K candidates of $\log p_{\text{base}}$ during decoding, achieving controllable text generation.
Different from the inference-time fusing methods and rely on additional outcome-based reward signal. In this paper, we propose to model $\log p_{\Delta}$ using a separate neutral network plugged during training, thereby efficiently achieving the transfer and enhancement of model capabilities.
@@ 3.1 Task Definition
Let $x$ denote an input sample. We utilize a pre-trained model $\pi_{\text{base}}({\bm{x}};\theta_{1})$ with fixed parameters $\theta_{1}$ that outputs logits ${\bm{z}}_{\text{base}}$:
${\bm{z}}_{\text{base}}(y_{t}|{\bm{x}},{\bm{y}}_{<t})=\pi_{\text{base}}({\bm{x}};\theta_{1}).$ (1)
To adapt this model to a new task without altering $\theta_{1}$, we introduce a delta model $\pi_{\Delta}(\,\circ\,;\theta_{2})$ with parameters $\theta_{2}$, producing adjustment logits:
@@
The final logits are computed by combining the pre-trained model’s logits with the delta model’s output:
@@
Let $p_{l}$ represent the label probabilities. Our objective is to minimize the Cross-Entropy loss function:
@@
By optimizing $\theta_{2}$, we aim to align $\log p_{\text{post}}$ with the target label distribution $p_{l}$, thereby enhancing performance on the new task while keeping the pre-trained parameters $\theta_{1}$ unchanged. Importantly, this process is equivalent to inversely learning an action-value function $Q(s,a)$ from demonstrations and $\pi_{\text{base}}$, where the states are $y_{<t}$ and the actions are $y_{t}$. In this context, actions at state $s$ receiving a higher ${\bm{z}}_{\text{post}}$ are encouraged by the value function based on the training demonstrations. Therefore we refer to this process as inverse value learning.
@@
    This framework enables both Weak-to-strong (Burns et al., 2024) or Strong-to-weak generations by introducing different amounts of parameters to parameterize $\pi_{\text{base}}$ and $\pi_{\Delta}$ (Mitchell et al., 2023). Since the delta model can be trained with small models and plugged into the larger models in the Weak-to-strong setting ($||\theta_{1}||>||\theta_{2}||$), the overall training time can be reduced compared to training a strong model directly.
@@
This formulation can introduce additional inference costs, as it requires obtaining logits from both the pre-trained and value models, resulting in extra computational overhead. Nevertheless, with more computational resources available and by adapting residual architecture, which will be discussed in Section 4.1, this will not introduce additional inference latency. Moreover, techniques like speculative decoding (Chen et al., 2023; Leviathan et al., 2023) could be employed to further optimize inference time. We leave this optimization to future work.
@@ 3.2 Linear Probes are Insufficient for Inverse Value Learning
Given the recent prominence of the surface alignment hypothesis (Zhou et al., 2023a; Lin et al., 2024), which posits that language models acquire the majority of their knowledge during pre-training. We have the reason to investigate whether a minimal set of parameters could sufficiently model the transformation between pre-trained and fine-tuned logits. To address this question, we conducted a preliminary experiment utilizing a single linear layer (see Appendix A.1 for details) to model $\pi_{\theta}$. We trained this model on ShareGPT (Chiang et al., 2023), a typical instruction-tuning dataset, and evaluated its efficacy.
Figure 2: MT-Bench scores of models with different sizes using linear probes for modeling $p_{\Delta}$ and full fine-tuning.
As shown in Figure 2, a single linear probe is insufficient to effectively model such a transformation. Specifically, while the model demonstrated proficiency in emulating the assistant’s style, it failed to capture and retrieve the task-specific knowledge required for problem-solving. We further categorized typical errors, as presented in Figure 5 in the Appendix. These results suggest that the relationship between pre-trained and fine-tuned logits is more complex than initially hypothesized, potentially involving a deeper neural network for more complex transformations or requiring a more sophisticated architectural modification.
Therefore, we further explored whether utilizing deeper networks could contribute to better modeling of this differential. Specifically, we employed a randomly initialized 7B Llama architecture as $\pi_{\Delta}$, subjected to the same training procedure to guide the alignment of an 1.1B pre-trained Llama Model. We observe that while it marginally outperforms the shallow linear probe in instruction-following capability, it still fell significantly short of full fine-tuning performance. This result indicates that employing pre-trained weights to predict the logits delta is necessary.
@@ 4.1 Connection Schemes
Figure 3: Cascade and Residual connection schemes for Inverse Value Learning. (a) Illustration of the two connection schemes. (b) MT-Bench scores for 1.1B models trained on the 1.1B base and generalized to the 7B base model using different connection schemes. “Cascade+” refers to the cascade value model where logits information from the pre-trained models is fused with the original token embeddings (marked as optional in Figure (a)) and used as input for the value model.
Based on the formulation in Section 3.1, we propose two distinct architectures to transform a pre-trained model into a value network and integrate it with a pre-trained model, namely Cascade and Residual, as illustrated in Figure 3 (a). It is important to note that both approaches maintain the internal structure of the LLM unchanged, focusing solely on the connection schemes between the pre-trained model and the value network.
@@
Alternatively, one may posit that the input derived from sampled text is not critical and the value model can effectively predict ${\bm{z}}_{\Delta}$ solely based on the logits from the pre-trained model. In this case:
@@
The full residual model takes previous tokens ${\bm{x}}$ as input and outputs the value scores:
${\bm{z}}_{\Delta}(y_{t}|{\bm{x}},{\bm{y}}_{<t})=\pi_{\Delta}({\bm{x}};\theta_{\text{residual}}).$
@@ 4.2 Plug-and-play transferability
Once a value model is trained on one backbone, it can be plugged into any other pre-trained models without further training, potentially achieving similar performance and behaviors as if it were post-trained. This capability stems from the feature of logits space steering (Mitchell et al., 2023). To further expand the boundaries of transferability, we investigate techniques to mitigate overfitting to the training-time backbone and methods for cross-vocabulary model transfer.
@@ 4.2.1 Regularization via Norm Constraint
Inevitably, the model tends to overfit the backbone it is trained with, thereby reducing its generalization capability on other pre-trained models. Applying appropriate regularization techniques during training can mitigate this issue. For example, we can directly constrain the sparsity of delta logits by incorporating an L1 norm term in the loss function. Specifically, this can be expressed as:
@@ 4.2.2 Vocabulary mapping
To adapt value-guided decoding to pre-trained and value models trained on different vocabularies and tokenizers, such as plugging a Llama-2-based value model to a Llama-3 family pre-trained model, a vocabulary mapping algorithm is essential. We adopt the MinED mapping algorithm  (Wan et al., 2024) for per-step mapping of the logits, which identifies mapping relationships between two sequences by minimizing their edit distance. Based on these mapping relationships, we can derive a vocabulary mapping matrix, which is then normalized and regularized to ensure the mapped logit values are not biased. During each decoding step, this mapping matrix is employed to transform the vocabulary space as stated in equation 3. The details of the algorithm and corresponding evaluation are shown in Appendix B.
@@ 5.3 Tasks
Table 1: Performance comparison of full fine-tuning and inverse value learning on instruction following, zero-shot and few-shot tasks. Note that value models of various parameter scales are trained together with base models of corresponding scales. For instance, “1.1B” indicates a 1.1B value model trained on a 1.1B backbone model. We further use a pink cell tag to denote a transfer setting. $\downarrow$” symbol indicates that lower values are better.
@@ 5.4 Results
Prior to conducting extensive transfer testing, we first evaluate which of the proposed schemes is more conducive to modeling ${\bm{z}}_{\Delta}$. Using an 1.1B model fine-tuned as the value model, we study the effects of the two structures proposed in Figure 2(a) on both the 1.1B backbone (used during training) and the 7B backbone (used for inference-time transfer). As shown in Figure 2(b), changes in the model’s connection scheme did not result in significant differences in general instruction-following ability, especially when transferring to other pre-trained models. Integrating logits from the base model slightly improves performance on the training backbone. However, when transfering to a larger backbone, there is a decrease in performance compared to the full residual structure. We attribute this to two potential factors: 1) Excessive input signals introduce noise to the cascade scheme, and 2) The value model in the cascade scheme no longer maintains a purely autoregressive form (i.e., it predicts the $p_{\Delta}$ based on the pre-trained logits at the same time step), widening the gap between pre-training and fine-tuning. Consequently, we adapt the full residual scheme for subsequent experiments in Table 1, given its superior performance and efficiency.
@@
The inverse learned value model demonstrated broad generalizability across models of different parameter scales, consistent with the findings of Mitchell et al. (2023). When an 1.1B parameter value model, trained on an 1.1B parameter backbone, is transferred to a 7B parameter base model, it retains approximately 80% of its instruction-following capability compared to full finetuning. Additionally, when a 7B parameter value model, jointly trained with a 7B base model, was transferred to a 13B parameter model, the instruction-following performance matched that of full finetuning. This transferability can similarly be scaled up to 70B.

### v2

@@ Abstract
As post-training processes utilize increasingly large datasets and base models continue to grow in size, the computational demands and implementation challenges of existing algorithms are escalating significantly. In this paper, we propose modeling the changes at the logits level during post-training using a separate neural network (i.e., the value network). After training this network on a small base model using demonstrations, this network can be seamlessly integrated with other pre-trained models during inference, enables them to achieve similar capability enhancements. We systematically investigate the best practices for this paradigm in terms of pre-training weights and connection schemes. We demonstrate that the resulting value network has broad transferability across pre-trained models of different parameter sizes within the same family, models undergoing continuous pre-training within the same family, and models with different vocabularies across families. In certain cases, it can achieve performance comparable to full-parameter fine-tuning. Furthermore, we explore methods to enhance the transferability of the value model and prevent overfitting to the base model used during training. [^1] Our code is open-source at https://github.com/icip-cas/inverse-value-learning
@@
Xinyu Lu1,2, Xueru Wen1,2 Yaojie Lu1, Bowen Yu3, Hongyu Lin1, Haiyang Yu3, Le Sun1,thanks:   Corresponding authors., Xianpei Han1, Yongbin Li3,11footnotemark: 1 1Chinese Information Processing Laboratory Institute of Software, Chinese Academy of Sciences 2University of Chinese Academy of Sciences 3Alibaba Group {luxinyu2021,wenxueru2022,luyaojie,hongyu,sunle,xianpei}@iscas.ac.cn {yubowen.ybw,yifei.yhy,shuide.lyb}@alibaba-inc.com
@@ 1 Introduction
We systematically investigate practical implementations for this framework. Specifically, we first study the importance of pre-trained weights in modeling the residual logits. Then, we examine two schemes for integrating the value network with pre-trained models: Cascade and Residual connections. Our experiments reveal that the residual connection scheme, where the value network predicts the delta logits based solely on previous text inputs, demonstrates superior transferability and efficiency. To improve the generalization ability of the value network, we incorporate regularization techniques, such as norm constraints, to mitigate overfitting to the base model during training. Additionally, we introduce a vocabulary mapping algorithm to facilitate effective cross-vocabulary transfer.
We assess our approach on multiple datasets and tasks, including general instruction following, zero-shot capabilities, and few-shot learning scenarios. Our results show that the value network exhibits broad transferability across pre-trained models of varying parameter sizes within the same model family, models undergoing continual pre-training, and even across different model families. In certain cases, our framework reaches close performance with full-parameter finetuning. This highlights the potential of our method for efficient and practical application.
@@ 2 Problem Statement and Background
The soaring computational costs of full-parameter updates with growing model sizes have driven researchers to explore efficient alternatives like low-rank parameter updates (Hu et al., 2022; Dettmers et al., 2024) or representation interventions (Wu et al., 2024) while freezing pre-trained weights. However, inconsistencies between the parameter spaces and representation spaces across different models stand as a key challenge for transferable post-training. In contrast, the logits space has better sharing properties and can serve as a communication channel for model interaction, enabling the transfer of capabilities between models (Hinton, 1999). Formally, the impact of post-training on the logits of a pretrained model can be expressed as:
@@
where $\log p_{\Delta}$ represents the change in the logits space during post-training. A series of inference-time proxy tuning studies focus on exploring $\log p_{\Delta}$. Mitchell et al. (2023) and Liu et al. (2024) obtained $\log p_{\Delta}$ by taking the difference between an existing instruction-following model and its corresponding base pre-trained model, i.e., $\log p_{\Delta}(y_{t}|{\bm{x}},{\bm{y}}_{<t})=\log\dfrac{p_{\pi^{*}}(y_{t}|{\bm{x}},{\bm{y}}_{<t})}{p_{\text{base}}(y_{t}|{\bm{x}},{\bm{y}}_{<t})}$. This difference can be regarded as an advantage function (Mitchell et al., 2023), or an implicit reward signal (Rafailov et al., 2024).
Another line of research using logits-based steering focuses on controlling model outputs to satisfy predefined attributes (e.g., style, toxicity, length). A key characteristic of these works is their reliance on trajectory-level rewards, such as scores from toxicity or topic classifiers. These studies investigate the conversion of trajectory-based rewards into token-level dense guidance (Khanov et al., 2024; Krause et al., 2020; Yang and Klein, 2021; Liu et al., 2021) to achieve attribute-compliant text generation.
Compared to the logits arithmetic works, we provide a more practical training-based solution that reduces the number of models needed during inference from three to two. Furthermore, in contrast to controllable text generation methods that rely on predefined outcome-based reward signals, we learn the reward function implicitly by modeling $\log p_{\Delta}$ using a separate neural network trained with demonstrations, thus efficiently achieving the transfer of post-training capabilities.
@@ 3.1 Task Definition
Let ${\bm{x}}$ denotes an input text, ${\bm{y}}$ denotes the model generated tokens. We utilize a pre-trained model $\pi_{\text{base}}$ with fixed parameters $\theta_{1}$ that outputs logits ${\bm{z}}_{\text{base}}$:
${\bm{z}}_{\text{base}}(y_{t}|{\bm{x}},{\bm{y}}_{<t})=\pi_{\text{base}}({\bm{x}},{\bm{y}}_{<t};\theta_{1}).$ (1)
To adapt this model to a new task without altering $\theta_{1}$, we introduce a delta model $\pi_{\Delta}(\,\circ\,;\theta_{2})$ with parameters $\theta_{2}$, producing steering logits:
@@
The final logits are computed by combining the pre-trained model’s logits with the delta model’s:
@@
Let $p_{l}$ represent the label probabilities (e.g., one-hot next-token distribution in supervised fine-tuning). The training objective is to minimize the Cross-Entropy loss function:
@@
By optimizing $\theta_{2}$, we aim to align ${\bm{z}}_{\text{post}}$ with the target label distribution $p_{l}$, thereby enhancing performance on the new task while keeping the pre-trained parameters $\theta_{1}$ unchanged. Importantly, this process is equivalent to inversely learning an action-value function $Q(s,a)$ based on demonstrations and $\pi_{\text{base}}$, where the states are $y_{<t}$ and the actions are $y_{t}$. In this context, the value function learned from training demonstrations assigns higher ${\bm{z}}_{\Delta}$ scores to actions that align with the demonstrated behaviors at each state $s$. Therefore we refer to this process as inverse value learning.
@@
    This framework enables both Weak-to-Strong (Burns et al., 2024) and Strong-to-Weak generalizations by introducing different scales of models to parameterize $\pi_{\text{base}}$ and $\pi_{\Delta}$ (Mitchell et al., 2023). Since the value model can be trained with small base models and plugged into the larger ones in the Weak-to-Strong setting ($|\theta_{1}|>|\theta_{2}|$), the overall training time can be reduced compared to training a strong model directly.
@@
This formulation can introduce additional inference costs, as it requires obtaining logits from both the pre-trained and value models, resulting in extra computational overhead. Nevertheless, with slightly more computational resources and by adapting residual architecture, which will be discussed in Section 4.1, this will not introduce additional inference latency. Moreover, techniques like speculative decoding (Chen et al., 2023; Leviathan et al., 2023) could be employed to further optimize inference time. We leave these optimizations to future work.
@@ 3.2 Linear Probes are Insufficient for Inverse Value Learning
Given the recent prominence of the surface alignment hypothesis (Zhou et al., 2023a; Lin et al., 2024), which posits that language models acquire the majority of their knowledge during pre-training. We have the reason to investigate whether a minimal set of parameters could sufficiently model the transformation between pre-trained and fine-tuned logits. To address this question, we conducted a preliminary experiment utilizing a single linear layer (see Appendix A.1 for details) to model ${\bm{z}}_{\Delta}$. We trained this model on ShareGPT (Chiang et al., 2023), a typical instruction-tuning dataset, and evaluated its efficacy.
Figure 2: MT-Bench scores of models with different sizes using linear probes for modeling $z_{\Delta}$ and full fine-tuning.
As illustrated in Figure 2, we find that a single linear probe is insufficient to effectively model such a transformation (more than a 3-point drop in MT-Bench scores). While the combined model successfully mimics the assistant’s conversational style, it fails to effectively apply the task-specific knowledge needed for problem-solving (typical failure modes shown in Figure 5 in the Appendix). This finding suggest that the relationship between pre-trained and fine-tuned logits is more sophisticated than initially hypothesized.
We further investigate whether the performance gap could be bridged by increasing model capacity while maintaining random initialization. Specifically, we employed a randomly initialized 7B Llama architecture as $\pi_{\Delta}$, following the identical training procedure to guide the alignment of an 1.1B pre-trained Llama model. While this model marginally outperforms the shallow linear probe in instruction-following capabilities, it still significantly underperforms compared to full fine-tuning. These results suggest that leveraging pre-trained weights is crucial for effective delta logits prediction.
@@ 4.1 Connection Schemes
Figure 3: Cascade and Residual connection schemes for Inverse Value Learning. (a) Illustration of the two connection schemes. (b) MT-Bench scores for 1.1B models trained on the 1.1B base and generalized to the 7B base model using different connection schemes. “Cascade+” refers to the cascade value model where logits information from the pre-trained models is fused with the original token embeddings (marked as optional in Figure (a)) and used as input for the value model. The final unembedding layer maps the residual stream from hidden space to the logits space, equivalent to the Linear Layer in the original Transformer literature (Vaswani et al., 2017).
Based on the formulation and observation in Section 3, we design two distinct architectures to transform a pre-trained model into a value network and integrate it with another pre-trained model, namely Cascade and Residual, as illustrated in Figure 3(a). It is important to note that both approaches maintain the internal structure of the LLM unchanged, focusing solely on the connection schemes between the pre-trained model and the value network.
@@
An alternative hypothesis suggests that the textual embedding input may be redundant, as the value model could potentially achieve comparable performance by exclusively utilizing the logit information from the pre-trained model. In this case:
@@
The full residual model takes previous tokens ${\bm{x}}$ and ${\bm{y}}_{<t}$ as inputs and outputs the value scores:
${\bm{z}}_{\Delta}(y_{t}|{\bm{x}},{\bm{y}}_{<t})=\pi_{\Delta}({\bm{x}},{\bm{y}}_{<t};\theta_{\text{residual}}).$
@@ 4.2 Plug-and-play transferability
Once a value model is trained on one backbone, it can be plugged into other pre-trained models without further training, potentially achieving similar performance and behaviors as if it were post-trained. This capability stems from the feature of logits space steering (Mitchell et al., 2023). To further expand the boundaries of transferability, we investigate techniques to mitigate overfitting to the training-time backbone and methods for cross-vocabulary model transfer.
@@ 4.2.1 Regularization via Norm Constraint
Inevitably, the value model tends to overfit the backbone it is trained with, thereby reducing its generalization capability on other pre-trained models. Applying appropriate regularization techniques during training can mitigate this issue. For example, we can directly constrain the sparsity of delta logits by incorporating an L1 norm term in the loss function. Specifically, this can be expressed as:
@@ 4.2.2 Vocabulary Mapping
To adapt value-guided decoding to pre-trained and value models trained on different vocabularies and tokenizers, such as plugging a Llama-2-based value model to a Llama-3 family pre-trained model, a vocabulary mapping algorithm is essential. We adopt the MinED mapping algorithm  (Wan et al., 2024) for per-step mapping of the logits, which identifies mapping relationships between two sequences by minimizing their edit distance. Based on these mapping relationships, we can derive a vocabulary mapping matrix, which is then normalized and regularized to ensure the mapped logit values are not biased. During each decoding step, this mapping matrix is employed to transform the vocabulary space as follows:
${\bm{z}}_{\text{post}}(y_{t}|{\bm{x}},{\bm{y}}_{<t})$ $={\bm{z}}_{\text{base}}(y_{t}|{\bm{x}},{\bm{y}}_{<t})\mathbf{W}$
$+{\bm{z}}_{\Delta}(y_{t}|{\bm{x}},{\bm{y}}_{<t}).$ (5)
where ${\bm{W}}\in\mathbb{R}^{n\times m}$ is the obtained transition matrix to map the size-$n$ pre-trained vocabulary to the size-$m$ vocabulary of the value model. The details of the algorithm and corresponding evaluation are shown in Appendix B.
@@ 5.3 Tasks
Table 1: Performance comparison of full fine-tuning and inverse value learning on instruction following, zero-shot and few-shot tasks. Note that value models of various parameter scales are trained together with base models of corresponding scales. For instance, “1.1B” indicates a 1.1B value model trained on a 1.1B backbone model. We further use a blue cell tag to denote a transfer setting. $\downarrow$” symbol indicates that lower values are better.
@@ 5.4 Results
Prior to conducting extensive transfer testing, we first evaluate which of the proposed connection schemes is more conducive to modeling ${\bm{z}}_{\Delta}$. Using an 1.1B model fine-tuned as the value model, we study the effects of the two schemes in Figure 3(a) on both the 1.1B backbone (used during training) and the 7B backbone (used for inference-time transfer). As shown in Figure 3(b), for the “Cascade+” and “Residual” schemes, difference in connection schemes did not result in significant variations in general instruction-following ability, especially when transferring to other pre-trained models. However, the token embedding serve as essential input features for the value model’s functionality (“Cascade” v.s. “Cascade+”). Meanwhile, integrating logits from the base model slightly improves performance on the training backbone. However, when transferring to a larger backbone, there is a decrease in performance compared to the full residual structure. We attribute this to two potential factors: 1) Excessive input signals introduce noise to the cascade scheme, and 2) The value model in the cascade scheme no longer maintains a purely autoregressive form (i.e., it predicts the ${\bm{z}}_{\Delta}$ based on the pre-trained logits at the same time step), widening the gap between pre-training and fine-tuning. Consequently, we adapt the full residual scheme for subsequent experiments in Table 1, given its superior performance and efficiency.
@@
The inverse learned value model demonstrated broad generalizability across models of different parameter scales, consistent with the findings of Mitchell et al. (2023). When an 1.1B parameter value model, trained on an 1.1B parameter backbone, is transferred to a 7B parameter base model, it retains large portion of its instruction-following capability compared to full finetuning. Additionally, when a 7B parameter value model, jointly trained with a 7B base model, was transferred to a 13B parameter model, the instruction-following performance matched that of full finetuning. This transferability can similarly be scaled up to 70B.